OpenStudy (anonymous):

Help me! y= -x^2-x+2 This involves graphing

OpenStudy (anonymous):

The graph should be a parabola. Use the quadratic formula \[ x_{1/2} = \frac{ -b \pm \sqrt{b^2 - 4ac}}{2a} \]To find the x intercepts of the graph. Then, find the y intercept by zeroing the x. Because it's -x^2, it should be concave downwards.

OpenStudy (anonymous):

Also, use the vertex formula: \[x_{vertex} = \frac{-b}{2a} \]The graph should look something like this: http://www4a.wolframalpha.com/Calculate/MSP/MSP23111a18eec318di138e00005d8ibg61baf62bbi?MSPStoreType=image/gif&s=31&w=299&h=130&cdf=RangeControl

OpenStudy (anonymous):

It factors as y=(x+2)(1-x)